OpenStudy (anonymous):

Hi everyone! Can someone discuss with me the purpose of a "total differential" and what it really represents? Thanks! :o)

OpenStudy (anonymous):

what do you think it is.

OpenStudy (anonymous):

well...the derivative of a function is the rate of change or the slope of that function at a given point y with respect to x...i get that part... so like if I have an equation y=x^2 + 2 the derivative is simply 2x...but what confuses me is that when you get an equation like f(x,y)=x^2y^2...if I take the derivative with respect to x, it's a bit confusing because there is a "y" in the function whereas in the first equation there is only an x to deal with...can you help me understand what I am finding the slope of? Is it a 3rd dimension or something?

OpenStudy (anonymous):

ok that messed up hold on

OpenStudy (anonymous):

omg...in your answer i see a bunch of black diamonds with question marks inside them...i don't think i am displaying what you wrote correctly

OpenStudy (anonymous):

try that sorry

OpenStudy (anonymous):

try what? the wiki link?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yeah that doesn't help, sorry...i need to discuss it, not read all the mathese...i appreciate the help but I could have looked that up on my own, which I did earlier...can you just try and explain it to me?

OpenStudy (anonymous):

was i correct about the total differential being the slope of both x and y in the 3rd dimension(z axis)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you know the mean part

OpenStudy (anonymous):

do you know df dimensions

OpenStudy (anonymous):

actually...

OpenStudy (anonymous):

my next question was if you could help me figure out the mechanics of solving a total differential...for example: given f(x,y)=x^2y^2 can you show me mechanically why the answer is df = 2xy^2 dx + 2x^2y dy I can't figure out how they get the dx and dy and df in their places

OpenStudy (anonymous):

@Kainui @jim_thompson5910

OpenStudy (anonymous):

I have looked at youtube but I can't seem to find a good video...can anyone refer me to a good video that explains the dx dy df thingy?

OpenStudy (anonymous):

Thanks Ganeshie...I will look right now

OpenStudy (anonymous):

quick question Ganeishi...

OpenStudy (anonymous):

so the video just simply says that the differential of f(x,y)=df as far as the left side of the equation goes... typically when you take the derivative of say f(x), you would write: d(f)/dx...since you are taking the total derivative of say f(x,y), is "df" like writing d(f)/dxdy or something like that? I am trying to understand where the "df" comes from mechanically

OpenStudy (anonymous):

d(f)/dxdy doesn't really make sense though because you would because it would look like df= stuff dxdy + morestuff dxdy rather than df=stuff dx + morestuff dy...can anyone explain what I am missing?

OpenStudy (anonymous):

or did I not watch enough of the video :o~ ?

OpenStudy (anonymous):

okay...I think I understand. "df" is not a number, so I can't expect to manipulate d(f(x,y)) like you normally would..."df" is "equal" to the "idea" of combining d(f)/dx +d(f)/dy +d(f)/dz = df :o) right?

OpenStudy (anonymous):

thanks!